# Periodic classical trajectories and quantum scars in many-spin systems

### v1

@@ II.1 General aspects
When it comes to the transition from classical to quantum thermalisation, one of the primary challenges here is the fact that the classical description is based on the notions of ergodicity and chaos of phase space trajectories, while quantum-mechanically the very notion of a phase trajectory is not definable – consequence of the Heisenberg uncertainty principle not allowing one to simultaneously determine the coordinates and momenta of particles with infinite precision. On the quantum side, chaotic instabilities and ergodicity become detached from each other. The common indicators of “quantum chaos”, such as the statistics of energy level spacings [51, 52] or eigenstate thermalisation [53, 54, 55, 56, 57, 58], involve the global properties of many-body energy eigenstates, and therefore, characterise their ergodicity. At the same time, recent research on out-of-time-ordered correlators (OTOCs) indicates[59, 60, 61, 62, 63] that chaotic instabilities in response to small deviations of initial conditions are, strictly speaking, absent in quantum systems with bounded local operators (e.g. spins 1/2) and finite-range interactions even in the thermodynamic limit. (Here, we sidestep the models with infinite-range interactions, where chaos and thermalisation properties may be rather special[64, 65, 66, 67].) Systems with finite-range interactions, however, may still exhibit a transient exponential OTOC growth representative of the classical chaotic limit, in particular, for lattices of spins $S>1/2$ [60, 62]. A similar condition, namely, $S\geq 3/2$, to observe the effect of classical periodic trajectories on quantum dynamics is reported in Section VI of the present work. The readers are referred to Refs.[68, 69, 70, 71, 72, 73] for various complementary aspects of the correspondence between classical and quantum spin dynamics.
Other efforts to identify the indicators and the mechanisms of the proper quantum thermalisation targeted the properties of the Hamiltonian matrix elements[74, 75], the role of quantum typicality[76, 77, 78, 79, 80, 81], generic features of the relaxation behavior[82, 83, 84, 85, 86], the frequency spectra of the observables[87, 88], quantum scrambling[89, 90, 91, 92, 93, 94, 95] and, more recently, the universal operator growth[88, 96, 97, 98]. All the above properties are subject to caution due to possible generic existence of exceptional quantum-scar states associated with the periodic classical orbits of the kind considered in this work.
@@ II.2 Quantum scars
Let us recall here that the concept of quantum scars as such was initially introduced by Heller[111] in the eighties for one particle moving in a chaotic two-dimensional billiard: it referred to atypical quantum eigenstates corresponding to classical periodic trajectories with short periods. When a classical periodic trajectory is sufficiently short, then the quasiclassical limit of the quantum dynamics implies the onset of quantum interference corrections associated with the trajectory closing onto itself before the chaotic Lyapunov instability suppresses such interference, and this corrections, in turn, amplify the probability of the corresponding eigenstate to remain localized around the classical trajectory as opposed to being evenly spread over the entire system which the typical eigenstates do. Numerous examples of low-dimensional quantum scars have been investigated since then – see, e.g., [112, 113]. One missing link between the original low-dimensional quantum scars and the recently discovered quantum many-body scars was that the latter were not connected to periodic classical trajectories. In other words, the quantum many-body scars as discovered were rather “distant relatives” to the quantum scars known from the physics of billiards. This issue of missing periodic trajectories was addressed in [114], although not in the classical limit but rather in the parameter space of quantum matrix product states.
@@ IV.5.1 Overview
We conjecture that the oscillatory dependence of $\lambda_{\text{p}}$ on $L$ in Fig.4, originates from the fact that, for a chain of a given length $L$, the allowed values of the wave number $q=\frac{2\pi}{L}n$, where $n=0,1,...,L-1$, cannot exactly match a certain value $q_{0}$ that is optimal for maximising the Lyapunov exponent $\lambda_{\text{p}}$. Let us denote that maximal value as $\lambda_{\text{p,max}}$. As the length of the chain increases, the allowed values of $q$ come increasingly closer to $q_{0}$, and, as a result, the oscillatory part of $\lambda_{\text{p}}(L)$ becomes gradually suppressed. We are dealing with a supposedly small difference between $q_{0}$ and the nearest allowed value $q_{\text{p}}=\frac{2\pi}{L}\ \text{round}\left(\frac{q_{0}L}{2\pi}\right)$, where function “$\text{round}(...)$” rounds its argument to the nearest integer value. It is thus reasonable to parameterise the departure from $\lambda_{\text{p,max}}$ by a quadratic dependence on $(q_{0}-q_{\text{p}})$:
@@ V.7 Mechanism C: Arnold diffusion
As mentioned earlier, we have encountered several examples indicated in Fig. 4, where neither Mechanism A nor Mechanism B are operational and hence the lifetime of the nearly quasiperiodic regime becomes anomalously long. We now put forward a heuristic case that the Mechanism C behind the destruction of the quasiperiodic regime in those examples is the Arnold diffusion.
The idea of Arnold diffusion goes back to Arnold’s article of 1964 [42], where an example was given, which implied that a stochastic regime of weakly nonintegrable dynamical systems with phase space dimensions larger than 4 should be unconstrained. The time required to explore the stochastic part of the energy shell was, however, exponentially long - proportional to the exponent of a negative power of the perturbation strength - and thus not accessible by a perturbation theory. The above proposition was contrasted with the phase spaces of dimensions $\leq 4$, where the KAM tori of the quasiperiodic motion surviving a finite small perturbation would separate the existing stochastic regions, thereby making the stochastic part of the energy shell disconnected. We note that the term “Arnold diffusion” does not mean a diffusion in a conventional sense but rather stands for unboundedness of the stochastic motion. Although Arnold diffusion has remained in the focus of active mathematical and physical research [125, 126, 127, 128, 129, 43, 130, 131, 49], the relevance of this phenomenon to realistic many-body physical systems has remained rather enigmatic.
@@
Finally, we note that the Arnold diffusion regime is similar to a variety of phenomena, where the time-translation symmetry is spontaneously broken: the relevant concepts discussed in the literature are time crystals [132, 133, 134, 135, 136, 137, 138] and time quasi-crystals [139, 140]. The present result adds a new member to this variety. Here we are dealing with the time translational symmetry of the original periodic trajectory that gets spontaneously broken by the Lyapunov instability into a regime, which may be called “quenched time quasi-crystall”, implying that there are no long-range correlations in the symmetry-broken state, and yet locally it looks like a time quasicrystal. The distinctive feature of our setting is that it involves neither periodic driving nor dissipation.
@@ Appendix D Functional form of $\lambda_{\text{p}}(J)$ near the separatrix $J=J^{*}$
where $T$ is the period of the trajectory. We further note that the separatrix has a symmetric 8-shaped form with the unstable fixed point in the middle. The spin orientation corresponding to the fixed point is to be denoted as $\mathbf{S}^{*}$. (Note: $\mathbf{S}^{*}$ is a three-dimensional vector for the one-spin representation of a periodic trajectory, while $\mathcal{S}^{*}$ is the corresponding point in the many-spin phase space.) The rotations near the separatrix cover half of that 8-shape, while the librations cover the entire 8-shape. Since the 8-shape is symmetric, it is sufficient in the case of librations to average the stretching rates in Eq.(32) only over the half of the libration period. Furthermore, since both the half of the libration trajectory and the rotation trajectory nearly coincide in the vicinity of the separatrix, the function $\lambda_{\text{p}}(J)$ tends to be largely symmetric with respect to small deviations from the separatrix $\Delta J\equiv J-J^{*}$ in agreement with the the visual impression from Figs. 7(a) and with the analysis below. Yet, this analysis also reveals a subdominant asymmetry of $\lambda_{\text{p}}(J)$.
@@
The time interval $T_{S}$ of the Lyapunov growth of $|\Delta\mathbf{S}(t)|$ is then followed by the time interval $T_{0}$ of large deviations from the fixed point; that interval is approximately independent of $|\Delta J|$. At the end of $T_{0}$, $|\Delta\mathbf{S}(t)|$ crosses again the threshold value $|\Delta\mathbf{S}|_{\lambda}$ and starts approaching the fixed point in the Lyapunov regime controlled by the negative Lyapunov exponent $-\lambda_{S}$. Due to the time-reversal symmetry of the Hamiltonian dynamics, it takes another time $T_{S}$ until the the trajectory arrives to the point of the closest approach to the fixed point, thereby completing the the period $T$. The period $T$ can now be represented as:

### v2

@@ II.1 General aspects
When it comes to the transition from classical to quantum thermalisation, one of the primary challenges here is the fact that the classical description is based on the notions of ergodicity and chaos of phase space trajectories, while quantum-mechanically the very notion of a phase trajectory is not definable – consequence of the Heisenberg uncertainty principle not allowing one to simultaneously determine the coordinates and momenta of particles with infinite precision. On the quantum side, chaotic instabilities and ergodicity become detached from each other. The common indicators of “quantum chaos”, such as the statistics of energy level spacings [51, 52] or eigenstate thermalisation [53, 54, 55, 56, 57, 58], involve the global properties of many-body energy eigenstates, and therefore, characterise their ergodicity. At the same time, recent research on out-of-time-ordered correlators (OTOCs) indicates[59, 60, 61, 62, 63] that chaotic instabilities in response to small deviations of initial conditions are, strictly speaking, absent in quantum systems with bounded local operators (e.g. spins 1/2) and finite-range interactions even in the thermodynamic limit. (Here, we sidestep the models with infinite-range interactions, where chaos and thermalisation properties may be rather special[64, 65, 66, 67].) Systems with finite-range interactions, however, may still exhibit a transient exponential OTOC growth representative of the classical chaotic limit, in particular, for lattices of spins $S>1/2$ [60, 62]. A similar condition, namely, $S\geq 3/2$, to observe the effect of classical periodic trajectories on quantum dynamics is reported in Section VI of the present work. The readers are referred to Refs.[68, 69, 70, 71, 72, 73, 74] for various complementary aspects of the correspondence between classical and quantum spin dynamics.
Other efforts to identify the indicators and the mechanisms of the proper quantum thermalisation targeted the properties of the Hamiltonian matrix elements[75, 76], the role of quantum typicality[77, 78, 79, 80, 81, 82], generic features of the relaxation behavior[83, 84, 85, 86, 87], the frequency spectra of the observables[88, 89], adiabatic eigenstate deformations[90, 91], quantum scrambling[92, 93, 94, 95, 96, 97, 98] and, more recently, the universal operator growth[89, 99, 100, 101]. All the above properties are subject to caution due to possible generic existence of exceptional quantum-scar states associated with the periodic classical orbits of the kind considered in this work.
@@ II.2 Quantum scars
Let us recall here that the concept of quantum scars as such was initially introduced by Heller[114] in the eighties for one particle moving in a chaotic two-dimensional billiard: it referred to atypical quantum eigenstates corresponding to classical periodic trajectories with short periods. When a classical periodic trajectory is sufficiently short, then the quasiclassical limit of the quantum dynamics implies the onset of quantum interference corrections associated with the trajectory closing onto itself before the chaotic Lyapunov instability suppresses such interference, and this corrections, in turn, amplify the probability of the corresponding eigenstate to remain localized around the classical trajectory as opposed to being evenly spread over the entire system which the typical eigenstates do. Numerous examples of low-dimensional quantum scars have been investigated since then – see, e.g., [115, 116]. One missing link between the original low-dimensional quantum scars and the recently discovered quantum many-body scars was that the latter were not connected to periodic classical trajectories. In other words, the quantum many-body scars as discovered were rather “distant relatives” to the quantum scars known from the physics of billiards. This issue of missing periodic trajectories was addressed in [117, 118], although not in the classical limit but rather in the parameter space of quantum matrix product states.
@@ IV.5.1 Overview
We conjecture that the oscillatory dependence of $\lambda_{\text{p}}$ on $L$ in Fig.4, originates from the fact that, for a chain of a given length $L$, the allowed values of the wave number $q=\frac{2\pi}{L}n$, where $n=0,1,...,L-1$, cannot exactly match a certain value $q_{0}$ that is optimal for maximising the Lyapunov exponent $\lambda_{\text{p}}$ in a chain of infinite length. Let us denote that maximal value as $\lambda_{\text{p,max}}$. As the length of the chain increases, the allowed values of $q$ come increasingly closer to $q_{0}$, and, as a result, the oscillatory part of $\lambda_{\text{p}}(L)$ becomes gradually suppressed. We are dealing with a supposedly small difference between $q_{0}$ and the nearest allowed value $q_{\text{p}}=\frac{2\pi}{L}\ \text{round}\left(\frac{q_{0}L}{2\pi}\right)$, where function “$\text{round}(...)$” rounds its argument to the nearest integer value. It is thus reasonable to parameterise the departure from $\lambda_{\text{p,max}}$ by a quadratic dependence on $(q_{0}-q_{\text{p}})$:
@@ V.7 Mechanism C: Arnold diffusion
As mentioned earlier, we have encountered several examples indicated in Fig. 4, where neither Mechanism A nor Mechanism B are operational and hence the lifetime of the nearly quasiperiodic regime becomes anomalously long. We now put forward a heuristic case that the Mechanism C behind the destruction of the nearly quasiperiodic regime in those examples is the Arnold diffusion. It requires that there are no other finite-time Lyapunov instabilities in the vicinity of the nearly quasiperiodic trajectory besides the one that initiated that trajectory. [The term “finite-time Lyapunov instability” implies an instability averaged over a time period much shorter than the very long duration of the Arnold diffusion regime – to be distinguished from the definition (9).]
The idea of Arnold diffusion goes back to Arnold’s article of 1964 [42], where an example was given, which implied that a stochastic regime of weakly nonintegrable dynamical systems with phase space dimensions larger than 4 should be unconstrained. The time required to explore the stochastic part of the energy shell was, however, exponentially long - proportional to the exponent of a negative power of the perturbation strength - and thus not accessible by a perturbation theory. The above proposition was contrasted with the phase spaces of dimensions $\leq 4$, where the KAM tori of the quasiperiodic motion surviving a finite small perturbation would separate the existing stochastic regions, thereby making the stochastic part of the energy shell disconnected. We note that the term “Arnold diffusion” does not mean a diffusion in a conventional sense but rather stands for unboundedness of the stochastic motion. Although Arnold diffusion has remained in the focus of active mathematical and physical research [129, 130, 131, 132, 133, 43, 134, 135, 136, 137, 138, 139, 49], the relevance of this phenomenon to realistic many-body physical systems has remained rather enigmatic.
@@
Finally, we note that the Arnold diffusion regime is similar to a variety of phenomena, where the time-translation symmetry is spontaneously broken: the relevant concepts discussed in the literature are time crystals [140, 141, 142, 143, 144, 145, 146, 147] and time quasi-crystals [148, 145, 149]. The present result adds a new member to this variety. Here we are dealing with the time translational symmetry of the original periodic trajectory that gets spontaneously broken by the Lyapunov instability into a regime, which may be called “quenched time quasi-crystal”, implying that there are no long-range correlations in the symmetry-broken state, and yet locally it looks like a time quasicrystal. The distinctive feature of our setting is that it involves neither periodic driving nor dissipation.
@@ VI.2 Quantum scars in finite-size spin clusters
The tendency of larger quantum spins $S$ to exhibit more pronounced evidence of quantum scarring can be understood as follows. For a billiard-like quantum many-body scar to be visible, the quantum dynamics should be close enough to a semiclassical limit, in which case, the difference between a quantum scar and a generic eigenstate is as dramatic as the difference between short periodic and ergodic classical trajectories. Without proximity to the semiclassical limit, fluctuations associated with strong quantum interference effects are, presumably, making it difficult to distinguish generic from nongeneric eigenstates. In our case, Fig. 16 illustrates that the accuracy of the semiclassical approximation for quantum spin dynamics indeed improves quickly with the growth of $S$.
@@ VII Summary and discussion
Note added: After submitting the present work for publication, we were alerted by the authors of Refs.[157, 158, 159] about their closely related publications on the relations between classical periodic orbits and quantum many-body scars.
@@ Appendix D Functional form of $\lambda_{\text{p}}(J)$ near the separatrix $J=J^{*}$
where $T$ is the period of the trajectory. We further note that the separatrix has a symmetric 8-shaped form with the unstable fixed point in the middle. The spin orientation corresponding to the fixed point is to be denoted as $\mathbf{S}^{*}$. (Note: $\mathbf{S}^{*}$ is a three-dimensional vector for the one-spin representation of a periodic trajectory, while $\mathcal{S}^{*}$ is the corresponding point in the many-spin phase space.) The rotations near the separatrix cover half of that 8-shape, while the librations cover the entire 8-shape. Since the 8-shape is symmetric, it is sufficient in the case of librations to average the stretching rates in Eq.(32) only over the half of the libration period. Furthermore, since both the half of the libration trajectory and the rotation trajectory nearly coincide in the vicinity of the separatrix, the function $\lambda_{\text{p}}(J)$ tends to be largely symmetric with respect to small deviations from the separatrix $\Delta J\equiv J-J^{*}$ in agreement with the visual impression from Figs. 7(a) and with the analysis below. Yet, this analysis also reveals a subdominant asymmetry of $\lambda_{\text{p}}(J)$.
@@
The time interval $T_{S}$ of the Lyapunov growth of $|\Delta\mathbf{S}(t)|$ is then followed by the time interval $T_{0}$ of large deviations from the fixed point; that interval is approximately independent of $|\Delta J|$. At the end of $T_{0}$, $|\Delta\mathbf{S}(t)|$ crosses again the threshold value $|\Delta\mathbf{S}|_{\lambda}$ and starts approaching the fixed point in the Lyapunov regime controlled by the negative Lyapunov exponent $-\lambda_{S}$. Due to the time-reversal symmetry of the Hamiltonian dynamics, it takes another time $T_{S}$ until the trajectory arrives to the point of the closest approach to the fixed point, thereby completing the period $T$. The period $T$ can now be represented as: